# A multi-frequency spaceborne radar perspective of deep convection

### v1

@@ Abstract
Global numerical weather models are starting to resolve atmospheric moist convection which comes with a critical need for observational constraints. One avenue for such constraints is spaceborne radar which tend to operate at three wavelengths, Ku-, Ka- and W-band. Many studies of deep convection in the past have primarily leveraged Ku-band because it is less affected by attenuation and multiple scattering. However, future spaceborne radar missions might not contain a Ku-band radar and thus considering the view of convection from Ka-band or W-band compared to the Ku-band would be useful. This study examines a coincident dataset between the Global Precipitation Measurement (GPM) Mission and CloudSat as well as the entire GPM record to compare convective characteristics across various wavelengths within deep convection. We find that W-band reflectivity (Z) tends to maximize near the Ku-band defined echo-top while Ka-band often maximizes 4-5 km below. The height of the maximum Z above the melting level for W-band and Ka-band do not linearly relate to the Ku-band maximum. However, using the full GPM record the Ka-band 30 dBZ echo-tops can be linearly related to the Ku-band 40 dBZ echo-top with an $R^{2}$ of 0.62 and a root mean squared error of about 1 km. The spatial distribution of echo-tops from Ka-band corresponds well to the Ku-band echo-tops, highlighting regions of relatively large ice water path. This paper suggests that Ka-band only missions, like NASA’s Investigation for Convective Updrafts, should be able to characterize global convection in a similar manner to a Ku-band system.
@@ 1 Introduction
For many places around the world, direct observations of storm characteristics are difficult to obtain. Regions over oceans, complex terrain, and other remote areas without access to high temporal and spatial resolution ground-based remote sensing or in-situ observations are common across the world. Evaluating the representation of storms in global numerical weather predictions is therefore a major challenge. This is now becoming particularly problematic since global numerical weather prediction models have grid-spacings that are approaching convective allowing scales (e.g., 4 km: DYAMOND; Stevens et al., 2019). One solution to help observe storms globally is to use spaceborne cloud and precipitation radars (see Battaglia et al., 2020, for a review)
@@
Apart from EarthCARE, none of the previous spaceborne radars have Doppler capabilities, and except for CPR, are generally focused on measurement of precipitation size hydrometeors near the surface. To date, the current spaceborne radar record does not have any direct dynamic information to characterize the dynamical processes defining the intensity of storms which would be valuable for assessing the representation of atmospheric moist convection in Earth system models. Additionally, the potential value of Doppler radar data from space (e.g., EarthCARE) for this purpose remains an unresolved topic of research. As an alternative, convective proxies derived from radar reflectivity profiles, like a storm’s height at a given reflectivity threshold, have been derived. The general idea for reflectivity-only convective metrics is that the higher the echo-top height altitude of a specific reflectivity value, the more intense the convection. The reasoning is that a stronger updraft can loft more massive hydrometeors to higher altitudes, thus producing larger reflectivities at higher altitudes. The exact reflectivity value adopted for the echo-top calculation varies because of the different saturation values for each wavelength and the radars’ minimum sensitivity thus offering different insights about the lofting of hydrometeors on convection. Reflectivity values for convective proxies range from 0 and 10 dBZ for CPR (e.g., Stephens and Wood, 2007; Luo et al., 2008; Takahashi and Luo, 2014) and 30 to 40 dBZ for TRMM (e.g., Zipser et al., 2006; Houze Jr. et al., 2015) and GPM’s KuPR (e.g., Skofronick-Jackson et al., 2018). The broad result of the echo-top convective proxies between missions paints a consistent picture of stronger convection being generally located over land, with more frequent locations of intense convection over hotspots like Central Africa, North-Central Argentina and the central United States (Zipser et al., 2006; Houze Jr. et al., 2015; Takahashi and Luo, 2014). The differences in echo-top heights of the different radar systems, not yet researched in detail, also offers potential insights into processes associated with detraining ice in upper levels of storms.
@@
Past papers have compared the profiles of reflectivity measured by CloudSat, TRMM and GPM. Some papers compared sensors to one another (TRMM to CloudSat; Sindhu and Bhat, 2013), and others compared the profiles with ground-based radars (Fall et al., 2013), showing the strengths of each radar. CloudSat was designed to observe clouds, the formation of precipitation and as well as precipitation that goes undetected by GPM (e.g., Berg et al., 2010; Behrangi et al., 2016), whereas TRMM and GPM were designed to detect larger hydrometeors and more intense liquid precipitation rates. Beyond the measured reflectivity profiles, retrieved precipitation rates, both rain and snow, have been compared between the satellites and ground-based estimates. Initial snow estimates from GPM were substantially lower than those of CloudSat (Casella et al., 2017; Skofronick-Jackson et al., 2019; Mroz et al., 2021), but could be improved by adjustments to assumptions in the solid phase retrievals shown by Chase et al. (2020, 2021, 2022b). For liquid precipitation, Hayden and Liu (2018) suggested the use of CloudSat precipitation retrievals for the lower precipitation rates (approximately 1 $\mathrm{mm\ hr^{-1}}$) and GPM for the larger precipitation rates to help create a more complete picture of the global precipitation. Liang et al. (2024) also suggested a combined dataset between CloudSat and GPM’s KaPR to better capture the full profile of hydrometeors.
@@
Interpretation of radar echoes suffers from complicating factors, including effects of attenuation (e.g., Meneghini and Kozu, 1990; Lhermitte, 1990), multiple scattering (e.g., Battaglia and Panegrossi, 2020) and non-Rayleigh scattering that get exacerbated as the operating frequency of the radar increases. These challenges are more prevalent in convection, where vertical motions support the formation of larger hydrometeors. Although these difficulties are understood, there has been a dearth of both qualitative and quantitative comparisons of Ku-, Ka- and W-band reflectivity to assess these effects.
@@
The research reported in this study compares Ku-, Ka- and W-band radar profiles measured within deep convection. Given Ku-band is the least affected by attenuation, multiple scattering and non-Rayleigh scattering, the primary objective is to compare the Ka-band measured profiles to the Ku-band radar profiles and quantify how much of the Ku-band information content is captured. Ka-band is the primary evaluation because this paper intends to evaluate how the emerging Ka-band only platforms such as INCUS, and tomorrow.io will perform in deep convection. This study further compares the Ku- and Ka-band profiles to those obtained by the W-band CPR of CloudSat. The intent of this comparison is not to highlight the benefits of combining the information provided by these different frequencies, as this will be a topic of a future study, but rather to provide some insight, both qualitative and quantitative, on the attenuation, non-Rayleigh scattering, and multiple scattering experienced by the millimeter wavelength radars (W-, and Ka- band) within deep convection in contrast to the Ku-band reflectivity. The aims of the study are achieved both by analysis of coincident GPM and CloudSat observations complied in the Turk et al. (2021) database and by analysis of the entire Ku- and Ka- band reflectivity GPM record. The primary focus of the analyses is limited to echoes above the melting level and near the tops of intense deep storms where convective updrafts are often maximized (e.g., Varble et al., 2014).
@@ 2.1 Radar descriptions
The two radars used in this study are the CloudSat Cloud Profiling Radar (CPR; Stephens et al., 2002) and the Global Precipitation Measurement mission’s Dual-frequency Precipitation Radar (DPR; Hou et al., 2014). The CPR is a W-band (94 GHz; 3 mm) non-scanning radar (i.e., single beam) with a vertical resolution of 500 m, a horizontal resolution of about 1.4 km and a minimum sensitivity of about -30 dBZ early in the life of the radar to approximately -25 dBZ toward the end of the mission life. The CPR was launched in 2006 into a polar orbit with an inclination of 98.2 degrees as part of NASA’s A-train (Stephens et al., 2008). While the radar was initially planned to operate for 24 months, it succeeded in operating for 17 years and the mission formally ended in December 2023. In 2011, a spacecraft battery anomaly occurred which prevented CPR from collecting observations during night (i.e., no sunlight on the solar array). CPR collected data regularly in the daylight only mode until the end of the mission, with interruptions to the record occurring because of reaction wheel failures and the introduction of a solution designed to sustain the measurement record (e.g., Hallowell et al., 2022).
@@ 2.3 GPM-DPR dataset
To extend the multi-wavelength comparison of storms to a larger sample size, the entire GPM-DPR dataset, ranging from March 2014 to November 2023, is also analyzed. Both the 2A.DPR files and the 2B.CMB files are acquired to enable the use of both the raw radar measurements (2A.DPR) and the combined radar-radiometer retrievals of precipitation rate and water content (Grecu et al., 2016; Olson, 2022). Data are sub-sampled in the same way as the coincident dataset. After all subsampling, more than 39 million profiles are available for the analysis described in Section 3.2.
@@ 3.1 Case Studies: Coincident Dataset
Fig. 3 presents the mean profile across the leading line of convection (white line in Fig. 2b-e). For this case, CPR 0 and 10 dBZ echo-tops are 2 - 4 km taller than the echo-top heights from the KaPR and KuPR. In fact, CPR observes echo up to 15 km, while the KuPR and KaPR only have echo up to 13 and 11 km respectively, primarily because of the lack of sensitivity of GPM’s radars (compared to CPR). The maximum reflectivity for each wavelength happens at different parts of the profile as well. CPR peaks at 17 dBZ at about 10 km, while KaPR maximizes at 31 dBZ at 6 km, KuPR at 41 dBZ at 5.6 km.
@@ 3.1.3 Case Study Discussion
For both cases, the general structure of the core convection of the storms aloft ($>$ 5km ASL) is captured no matter the wavelength, radar resolution (i.e., along-track spacing) or sensitivity. The differences arise from the echo-top height locations, the anvil characteristics and the depth at which the shorter wavelengths become affected by attenuation and multiple scattering. Contextualizing these two cases of very strong convection in the broader convective proxy literature, both cases would be used in CPR studies of deep convective clouds with 10 dBZ echo-tops exceeding 10 km (e.g., Takahashi et al., 2023).Meanwhile from the KuPR perspective, the MCS on 12 Jun 2018 would be within the top $0.33\%$ of storms and the more isolated storm from 08 May 2019 would be within the top $0.03\%$ within the Zipser et al. (2006) framework (see Fig. 6 in Skofronick-Jackson et al., 2018). In the University of Washington classifications both storms are considered the Deep and Wide classification (Houze Jr. et al., 2007), but note the part sampled by the three radars for the MCS case was likely not the most intense part of the storm (stronger GMI brightness temperature deficits were found outside the DPR swath; not shown).
@@
These two cases provide insight on how to interpret single frequency missions when examining the strongest convective storms. It is known that shorter wavelength radars suffer from more attenuation (e.g., Table 4.7 in Meneghini and Kozu, 1990) and multiple scattering effects (e.g., Battaglia et al., 2010), which suggests missions that only have Ka- or W- band radar are likely to have complications with retrieving information within deep convection. Both examples discussed were intense convection and contained hydrometeors responsible for strong attenuation and multiple scattering. The combined radar-radiometer retrieval from GPM shows large amounts of ice mass and estimated the amount of two-way path integrated attenuation at 42 and 59 dB for Ka-band for the two cases respectively. These are likely extreme cases in anticipating the potential signal of the shorter wavelengths in deep convection. Despite the strong nature of these convective profiles, the Ka-band information matches well with the Ku-band for about 5 km of depth (i.e., echo-top to 6 km in the first case). This is encouraging for Ka-band only missions, like INCUS and tomorrow.io, because there is clear signal from the Ka-band in the upper portions of the storms where convective motions, and thus convective mass flux, are the greatest (e.g., Varble et al., 2014). The W-band has a similar signal but, as expected, the depth of profile where it mimics the Ku-band is limited to about the first 2 km (starting from the KuPR echo-top). In the next section, we use the entire coincident dataset to draw some statistical relationships between echo-top heights between sensors as well as the general profile characteristics.
@@ 3.2 Coincident Dataset Statistics
Using the full coincident dataset (approximately 5000 points) we first compare the height of the maximum of reflectivity above the melting level (MAML) between the shorter wavelengths (Ka- and W- band) and Ku-band (Fig. 6). It is not expected that the MAML for the shorter wavelengths should be identical to Ku-band because of the nuances of attenuation and non-Rayleigh scattering but the magnitude of the correlation could show that the shorter wavelengths have a similar signal in convective intensity. For the entire database considered, the MAML derived from both the Ka- and W- band show relatively weak correspondence to the Ku-band based heights (Fig. 6), with small coefficients of determination ($R^{2}$). Thus, the convective proxy of MAML for for KaPR and CPR are not reproduced in the same manner in KuPR data.
The correspondences between the radar profiles are measurably closer when expressed in terms of echo top heights (Fig. 7). $R^{2}$ values follow expectation with generally decreasing $R^{2}$ with larger differences in reflectivity thresholds. The largest $R^{2}$ values are between the 30 dBZ KaPR height and the 30 dBZ KuPR height, with a value of almost 0.9 (blue line Fig. 7), while the smallest $R^{2}$ is between CPR’s 0 dBZ echo-top height and KuPR’s 40 dBZ echo-top height with a value less than 0.2. Considering the top of the KuPR echo (heightStromTop), the CPR 10 dBZ echo-top corresponds reasonably well with an $R^{2}$ of 0.74. All linear fits between different echo-tops are shown in Table 1, along with the root mean squared error if the linear fit was used to translate KaPR or CPR echo-top height to a KuPR echo-top height. For example, if the KaPR 20 dBZ echo-top was translated to a KuPR 20dBZ echo-top, the root mean squared error is 0.62 km. Note that the linear regression coefficients in Table 1 for the KaPR to KuPR echo-tops are near 1 for both the 20 dBZ and 30 dBZ echtops, implying you can use the KaPR echo-top interchangeably with the KuPR echo-top.
@@
$\mathrm{KaPR_{MAML}}$ $\mathrm{KuPR_{MAML}}$ 0.28 0.82 0.52 0.96
$\mathrm{KaPR_{20}}$ $\mathrm{KuPR_{20}}$ 0.96 0.99 0.37 0.62
$\mathrm{KaPR_{30}}$ $\mathrm{KuPR_{30}}$ 0.87 0.98 0.37 1.03
$\mathrm{KaPR_{30}}$ $\mathrm{KuPR_{40}}$ 0.71 0.69 0.25 1.30
$\mathrm{CPR_{MAML}}$ $\mathrm{KuPR_{MAML}}$ 0.05 0.58 2.1 1.85
$\mathrm{CPR_{0}}$ heightStormTop 0.59 0.66 0.95 1.80
$\mathrm{CPR_{10}}$ heightStormTop 0.74 0.83 1.24 1.56

### v2

@@ Abstract
Global numerical weather models are starting to resolve atmospheric moist convection which comes with a critical need for observational constraints. One avenue for such constraints is spaceborne radar which tend to operate at three wavelengths, Ku-, Ka- and W-band. Many studies of deep convection in the past have primarily leveraged Ku-band because it is less affected by attenuation and multiple scattering. However, future spaceborne radar missions might not contain a Ku-band radar and thus considering the view of convection from Ka-band or W-band compared to the Ku-band would be useful. This study examines a coincident dataset between the Global Precipitation Measurement (GPM) Mission and CloudSat as well as the entire GPM record to compare convective characteristics across various wavelengths within deep convection. We find that W-band reflectivity (Z) tends to maximize near the Ku-band defined echo-top while Ka-band often maximizes 4-5 km below. The height of the maximum Z above the melting level for W-band does not linearly relate to the Ku-band maximum. However, using the full GPM record the Ka-band 30 dBZ echo-tops can be linearly related to the Ku-band 40 dBZ echo-top with an $R^{2}$ of 0.62 and a root mean squared error of about 1 km. The spatial distribution of echo-tops from Ka-band corresponds well to the Ku-band echo-tops, highlighting regions of relatively large ice water path. This paper suggests that Ka-band only missions, like NASA’s Investigation for Convective Updrafts, should be able to characterize global convection in a similar manner to a Ku-band system.
@@ 1 Introduction
For many places around the world, direct observations of storm characteristics are difficult to obtain. Regions over oceans, complex terrain, and other remote areas without access to high temporal and spatial resolution ground-based remote sensing or in-situ observations are common across the world. Evaluating the representation of storms in global numerical weather predictions is therefore a major challenge. This is now becoming particularly problematic since global numerical weather prediction models have grid-spacings that are approaching convective allowing scales (e.g., 4 km: DYAMOND; Stevens et al., 2019). One solution to help observe storms globally is to use spaceborne cloud and precipitation radars (see Battaglia et al., 2020, for a review).
@@
Apart from EarthCARE, none of the previous spaceborne radars have Doppler capabilities, and except for CPR, are generally focused on measurement of precipitation size hydrometeors near the surface. To date, the current spaceborne radar record does not have any direct dynamic information to characterize the dynamical processes defining the intensity of storms which would be valuable for assessing the representation of atmospheric moist convection in Earth system models. Additionally, the potential value of Doppler radar data from space (e.g., EarthCARE) for this purpose remains an unresolved topic of research, but has been investigated using radar forward simulators and numerical weather prediction (Kollias et al., 2022). As an alternative, convective proxies derived from radar reflectivity profiles, like a storm’s height at a given reflectivity threshold, have been derived. The general idea for reflectivity-only convective metrics is that the higher the echo-top height of a specific reflectivity value, the more intense the convection. The reasoning is that a stronger updraft can loft more massive hydrometeors to higher altitudes, thus producing larger reflectivities at higher altitudes. The exact reflectivity value adopted for the echo-top calculation varies because of the different saturation values for each wavelength and the radars’ minimum sensitivity thus offering different insights about the lofting of hydrometeors on convection. Reflectivity values for convective proxies range from 0 and 10 dBZ for CPR (e.g., Stephens and Wood, 2007; Luo et al., 2008; Takahashi and Luo, 2014) and 30 to 40 dBZ for TRMM (e.g., Zipser et al., 2006; Houze Jr. et al., 2015) and GPM’s KuPR (e.g., Skofronick-Jackson et al., 2018). The broad result of the echo-top convective proxies between missions paints a consistent picture of stronger convection being generally located over land, with more frequent locations of intense convection over hotspots like Central Africa, North-Central Argentina and the central United States (Zipser et al., 2006; Houze Jr. et al., 2015; Takahashi and Luo, 2014). The differences in echo-top heights of the different radar systems, not yet researched in detail, also offers potential insights into processes associated with detraining ice in upper levels of storms.
@@
Past papers have compared the profiles of reflectivity measured by CloudSat, TRMM and GPM. Some papers compared sensors to one another (TRMM to CloudSat; Sindhu and Bhat, 2013), and others compared the profiles with ground-based radars (Fall et al., 2013), showing the strengths of each radar. CloudSat was designed to observe clouds, the formation of precipitation and as well as precipitation that goes undetected by GPM (e.g., Berg et al., 2010; Behrangi et al., 2016), whereas TRMM and GPM were designed to detect larger hydrometeors and more intense liquid precipitation rates. Beyond the measured reflectivity profiles, retrieved precipitation rates, both rain and snow, have been compared between the satellites and ground-based estimates. Initial snow estimates from GPM were substantially lower than those of CloudSat (Casella et al., 2017; Skofronick-Jackson et al., 2019; Mroz et al., 2021), but could be improved by adjustments to assumptions in the solid phase retrievals shown by Chase et al. (2020, 2021, 2022). For liquid precipitation, Hayden and Liu (2018) suggested the use of CloudSat precipitation retrievals for the lower precipitation rates (approximately 1 $\mathrm{mm\ hr^{-1}}$) and GPM for the larger precipitation rates to help create a more complete picture of the global precipitation. Liang et al. (2024) also suggested a combined dataset between CloudSat and GPM’s KaPR to better capture the full profile of hydrometeors.
@@
Interpretation of radar echoes suffers from complicating factors, including effects of attenuation (e.g., Meneghini and Kozu, 1990; Lhermitte, 1990), multiple scattering (e.g., Battaglia et al., 2010) and non-Rayleigh scattering that get exacerbated as the operating frequency of the radar increases. These challenges are more prevalent in convection, where vertical motions support the formation of larger hydrometeors. Although these difficulties are understood, there has been a dearth of both qualitative and quantitative comparisons of Ku-, Ka- and W-band reflectivity to assess these effects.
@@
The research reported in this study compares Ku-, Ka- and W-band radar profiles measured within deep convection. Given Ku-band is the least affected by attenuation, multiple scattering and non-Rayleigh scattering, the primary objective is to compare the Ka-band measured profiles to the Ku-band radar profiles and quantify how much of the Ku-band information content is captured. Ka-band is the primary focus of this evaluation because this paper intends to evaluate how the emerging Ka-band only platforms, such as INCUS and tomorrow.io, will perform in deep convection. Furthermore, the choice of shorter wavelength radars (Ka- and W-) might be the preferred wavelength for future spaceborne missions because of the added cost of Ku-band systems and the difficulties in accommodating the requisite antenna sizes on smaller space platforms (i.e., Ku-band systems require a larger antenna to achieve a similar sensitivity to Ka-band and W-band). This study further compares the Ku- and Ka-band profiles to those obtained by the W-band CPR of CloudSat. The intent of this comparison is not to highlight the benefits of combining the information provided by these different frequencies, as this will be a topic of a future study, but rather to provide some insight, both qualitative and quantitative, on the attenuation, non-Rayleigh scattering, and multiple scattering experienced by the millimeter wavelength radars (W-, and Ka- band) within deep convection in contrast to the Ku-band reflectivity. The aims of the study are achieved both by analysis of coincident GPM and CloudSat observations complied in the Turk et al. (2021) database and by analysis of the entire Ku- and Ka- band reflectivity GPM record. The primary focus of the analyses is limited to echoes above the melting level and near the tops of intense deep storms where convective updrafts are often maximized (e.g., Varble et al., 2014).
@@ 2.1 Radar descriptions
The two radars used in this study are the CloudSat Cloud Profiling Radar (CPR; Stephens et al., 2002) and the Global Precipitation Measurement mission’s Dual-frequency Precipitation Radar (DPR; Hou et al., 2014). The CPR is a W-band (94 GHz; 3 mm) non-scanning radar (i.e., single beam) with a vertical resolution of 500 m, a horizontal resolution of about 1.4 km and a minimum sensitivity of about -30 dBZ early in the life of the radar to approximately -25 dBZ toward the end of the mission life. The CPR was launched in 2006 into a sun synchronous polar orbit with an inclination of 98.2 degrees as part of NASA’s A-train (Stephens et al., 2008). The approximate overpass times of CPR is 1:30 am/pm local time at the equator. While the radar was initially planned to operate for 24 months, it succeeded in operating for 17 years and the mission formally ended in December 2023. In 2011, a spacecraft battery anomaly occurred which prevented CPR from collecting observations during night (i.e., no sunlight on the solar array). CPR collected data regularly in the daylight only mode until the end of the mission, with interruptions to the record occurring because of reaction wheel failures and the introduction of a solution designed to sustain the measurement record (e.g., Hallowell et al., 2022).
@@ 2.3 GPM-DPR dataset
To extend the multi-wavelength comparison of storms to a larger sample size, the entire GPM-DPR dataset, ranging from March 2014 to November 2023, is also analyzed. Both the 2A.DPR files and the 2B.CMB files are acquired to enable the use of both the uncorrected measured radar reflectivities (2A.DPR) and the combined radar-radiometer retrievals of precipitation rate and water content (Grecu et al., 2016; Olson, 2022). Data are sub-sampled in the same way as the coincident dataset. After all subsampling, more than 39 million profiles are available for the analysis described in Section 3.2.
@@ 3.1 Case Studies: Coincident Dataset
Fig. 3 presents the mean profile across the leading line of convection (white line in Fig. 2b-e). For this case, CPR 0 and 10 dBZ echo-tops are 2 - 4 km taller than the echo-top heights from the KaPR and KuPR. In fact, CPR observes echo up to 15 km, while the KuPR and KaPR only have echo up to 13 and 11 km respectively, primarily because of the lack of sensitivity of GPM’s radars (compared to CPR). The maximum reflectivity for each wavelength occurs at different parts of the profile as well. CPR peaks at 17 dBZ at about 10 km, while KaPR maximizes at 31 dBZ at 6 km, KuPR at 41 dBZ at 5.6 km.
@@ 3.1.3 Case Study Discussion
For both cases, the general structure of the core convection of the storms aloft ($>$ 5km ASL) is captured no matter the wavelength, radar resolution (i.e., along-track spacing) or sensitivity. The differences arise from the echo-top height locations, the anvil characteristics and the depth at which the shorter wavelengths become affected by attenuation and multiple scattering. Contextualizing these two cases of very strong convection in the broader convective proxy literature, both cases would be used in CPR studies of deep convective clouds with 10 dBZ echo-tops exceeding 10 km (e.g., Takahashi et al., 2023). Meanwhile from the KuPR perspective, the MCS on 12 June 2018 would be within the top $0.33\%$ of storms and the more isolated storm from 08 May 2019 would be within the top $0.03\%$ within the Zipser et al. (2006) framework (see Fig. 6 in Skofronick-Jackson et al., 2018). In the University of Washington classifications both storms are considered the Deep and Wide classification (Houze Jr. et al., 2007), but note the part sampled by the three radars for the MCS case was likely not the most intense part of the storm (stronger GMI brightness temperature deficits were found outside the DPR swath; not shown).
@@
These two cases provide insight on how to interpret single frequency missions when examining the strongest convective storms. It is known that shorter wavelength radars suffer from more attenuation (e.g., Table 4.7 in Meneghini and Kozu, 1990) and multiple scattering effects (e.g., Battaglia et al., 2010), which suggests missions that only have Ka- or W- band radar are likely to have complications with retrieving information within deep convection. Both examples discussed were intense convection and contained hydrometeors responsible for strong attenuation and multiple scattering. The combined radar-radiometer retrieval from GPM shows large amounts of ice mass and estimated the amount of two-way path integrated attenuation at 42 and 59 dB for Ka-band for the two cases respectively. The goal of this paper is to assess the information contained in the signals of shorter wavelength radars in deep convection, and evaluate the associated attenuation, multiple scattering, and non-Rayleigh scattering. The cases presented here demonstrate all of these challenges and are within tail of the distribution of deep convective samples. Despite the strong nature of these convective profiles, the Ka-band information matches well with the Ku-band for about 5 km of depth (i.e., echo-top to 6 km in the first case). This is encouraging for Ka-band only missions, like INCUS and tomorrow.io, because there is clear signal from the Ka-band in the upper portions of the storms where convective motions, and thus convective mass flux, are the greatest (e.g., Varble et al., 2014). The W-band has a similar signal but, as expected, the depth of profile where it mimics the Ku-band is limited to about the first 2 km (starting from the KuPR echo-top). In the next section, we use the entire coincident dataset to draw some statistical relationships between echo-top heights between sensors as well as the general profile characteristics.
@@ 3.2 Coincident Dataset Statistics
Using the full coincident dataset (approximately 5000 points) we first compare the height of the maximum of reflectivity above the melting level (MAML) between the shorter wavelengths (Ka- and W- band) and Ku-band (Fig. 6). It is not expected that the MAML for the shorter wavelengths should be identical to Ku-band because of the nuances of attenuation and non-Rayleigh scattering but the magnitude of the correlation could show that the shorter wavelengths have a similar signal in convective intensity. For the entire database considered, the MAML derived from Ka-band compared to Ku-band shows a weak linear relationship ($R^{2}$: 0.47; Fig. 6a) while W-band and Ku-band have no linear relationship ($R^{2}$: 0.05; Fig. 6b).
The correspondences between the radar profiles are measurably closer when expressed in terms of echo top heights (Fig. 7). $R^{2}$ values follow expectation with generally decreasing $R^{2}$ with larger differences in reflectivity thresholds. The largest $R^{2}$ values are between the 30 dBZ KaPR height and the 30 dBZ KuPR height, with a value of 0.83 (blue line Fig. 7), while the smallest $R^{2}$ is between CPR’s 0 dBZ echo-top height and KuPR’s 40 dBZ echo-top height with a value less than 0.15. Considering the top of the KuPR echo (heightStormTop), the CPR 10 dBZ echo-top corresponds reasonably well with an $R^{2}$ of 0.74. All linear fits between different echo-tops are shown in Table 1, along with the root mean squared error if the linear fit was used to translate KaPR or CPR echo-top height to a KuPR echo-top height. For example, if the KaPR 20 dBZ echo-top was translated to a KuPR 20dBZ echo-top, the root mean squared error is 0.77 km. Note that the linear regression coefficients in Table 1 for the KaPR to KuPR echo-tops are near 1 for both the 20 dBZ and 30 dBZ echtops, implying you can use the KaPR echo-top interchangeably with the KuPR echo-top.
@@
$\mathrm{KaPR_{MAML}}$ $\mathrm{KuPR_{MAML}}$ 0.47 0.93 0.37 0.78
$\mathrm{KaPR_{20}}$ $\mathrm{KuPR_{20}}$ 0.89 0.90 1.12 0.77
$\mathrm{KaPR_{30}}$ $\mathrm{KuPR_{30}}$ 0.83 0.95 1.27 1.05
$\mathrm{KaPR_{30}}$ $\mathrm{KuPR_{40}}$ 0.59 0.60 0.77 1.31
$\mathrm{CPR_{MAML}}$ $\mathrm{KuPR_{MAML}}$ 0.05 0.54 1.96 1.83
$\mathrm{CPR_{0}}$ heightStormTop 0.51 0.55 3.28 1.67
$\mathrm{CPR_{10}}$ heightStormTop 0.59 0.65 3.36 1.56